# Strong nebular emissions associated with Mg ii absorptions detected in the SDSS spectra of background quasars

### v1

@@ 2 The GOTOQS sub-sample
![Refer to caption](x1.png)
![Refer to caption](x2.png)
![Refer to caption](x3.png)
@@ 3 Observations and data reductions
![Refer to caption](x4.png)
![Refer to caption](x5.png)
![Refer to caption](x6.png)
![Refer to caption](x7.png)
![Refer to caption](x8.png)
![Refer to caption](x9.png)
![Refer to caption](x10.png)
![Refer to caption](x11.png)
![Refer to caption](x12.png)
![Refer to caption](x13.png)
![Refer to caption](x14.png)
![Refer to caption](x15.png)
![Refer to caption](x16.png)
![Refer to caption](x17.png)
![Refer to caption](x18.png)
![Refer to caption](x19.png)
![Refer to caption](x20.png)
![Refer to caption](x21.png)
![Refer to caption](x22.png)
![Refer to caption](x23.png)
![Refer to caption](x24.png)
![Refer to caption](x25.png)
![Refer to caption](x26.png)
![Refer to caption](x27.png)
![Refer to caption](x28.png)
![Refer to caption](x29.png)
![Refer to caption](x30.png)
![Refer to caption](x31.png)
![Refer to caption](x32.png)
![Refer to caption](x33.png)
![Refer to caption](x34.png)
![Refer to caption](x35.png)
![Refer to caption](x36.png)
![Refer to caption](x37.png)
![Refer to caption](x38.png)
![Refer to caption](x39.png)
![Refer to caption](x40.png)
![Refer to caption](x41.png)
![Refer to caption](x42.png)
![Refer to caption](x43.png)
@@ 4.1 Measuring the impact parameters
![Refer to caption](x44.png)
![Refer to caption](x45.png)
![Refer to caption](x46.png)
![Refer to caption](x47.png)
@@ 4.3 Line of sight reddening towards GOTOQ
![Refer to caption](x51.png)
Assuming the Mg ii absorbing gas follows the dust extinction similar to that of the small magellanic clouds (SMC) (Gordon et al., 2003), we calculate the color excess of the background quasar spectra using a standard quasar spectral template (Selsing et al., 2016). We first redshift the quasar template to the redshift of the background quasar and then apply the SMC dust extinction law at the Mg ii absorption redshift to match it with the shape of the quasar spectrum in hand by varying the dust extinction coefficient, . The color excess, for our sample ranges from to 0.422, with a median value of 0.058. The measured values of are provided in the last column of Table 1. Given the median of 2.21 Å, the expected color excess towards this GOTOQ sub-sample is (Budzynski & Hewett, 2011). Therefore, the line of sight reddening towards the GOTOQ is significantly higher than the typical Mg ii absorption systems of similar . Wild & Hewett (2005) have measured the for a sample of Ca ii absorbers. Their values range from to 0.30 with a median value of 0.042.
@@ 4.4 Estimation of galaxy parameters
There is a correlation found between stellar mass and impact parameter for the general population of Mg ii absorbers Guha et al. (2022). It has also been found that for a given impact parameter, the estimated for the USMgii absorbers tend to be higher compared to the Mg ii absorbers with low . These correlations imply that (i) at a given impact parameter, the absorbers with large tend to have large and (ii) for a given galaxies at higher impact parameters tend to have higher values of . Therefore, if GOTOQs follow this trend, we expect the for the host galaxies to be typically less than what has been found for USMgii host galaxies at high impact parameters. Indeed, Guha & Srianand (2023) have measured over the redshift range . Here, we check whether such measurements for individual cases in the case of GOTOQs follow the above-mentioned trend.
Accurate photometric measurements are required in several bands to measure the stellar masses () of GOTOQs using SED fitting techniques. This is possible only for 12 of the 39 objects with clear photometric extensions in our sample. For these twelve systems, the angular separations between the quasars and the foregrounds are large, so reliable fluxes in different photometric bands can be measured by subtracting the quasar PSF. We computed the foreground galaxy fluxes in the grz bands of the DECaLS photometry. Then, we fitted the spectral energy distributions (SED) using the publicly available code BAGPIPES (Carnall et al., 2018). During the fit, we assumed a delayed star formation history (Carnall et al., 2019) and kept the redshift constant as obtained from the [O ii] emission line. Obtained stellar masses of the GOTOQ with photometric counterparts are given in column 3 of Table 4. In Figure 9, we show plot vs. D for these objects and also show the best fit obtained for the host galaxies of USMgii and that obtained for the MAGIICAT galaxies. However, it is important to note that the correlation observed for the GOTOQs may be influenced by the mass-size relation of galaxies (Ward et al., 2024). Since more massive galaxies tend to be larger, massive GOTOQs are also larger in size and, consequently, may have larger impact parameters. The predicted stellar masses are consistent within the allowed uncertainties in these cases. It will be good to measure the host galaxy properties for more GOTOQs using HST photometry or AO-assisted ground-based imaging. This will provide more insights into how stellar mass controls the equivalent width distributions at low impact parameters.

### v2

@@ 4.3 Line of sight reddening towards GOTOQ
Assuming the Mg ii absorbing gas follows the dust extinction similar to that of the small magellanic clouds (SMC) (Gordon et al., 2003), we calculate the color excess of the background quasar spectra using a standard quasar spectral template (Selsing et al., 2016). We first redshift the quasar template to the redshift of the background quasar and then apply the SMC dust extinction law at the Mg ii absorption redshift to match it with the shape of the quasar spectrum in hand by varying the dust extinction coefficient, . The color excess, for our sample ranges from to 0.422, with a median value of 0.058. The measured values of are provided in the last column of Table 1. However, it is important to note that the highest line of sight reddening is observed for the system J1314+0330. As previously mentioned, this system is incorrectly identified as a GOTOQ. Therefore, like before, we do not include it in the analysis presented here. Given the median of 2.21 Å, the expected color excess towards this GOTOQ sub-sample is (Budzynski & Hewett, 2011). Therefore, the line of sight reddening towards the GOTOQ is significantly higher than the typical Mg ii absorption systems of similar . Wild & Hewett (2005) have measured the for a sample of Ca ii absorbers. Their values range from to 0.30 with a median value of 0.042.
@@ 4.4 Estimation of galaxy parameters
There is a correlation found between stellar mass and impact parameter for the general population of Mg ii absorbers in Guha et al. (2022). It has also been found that for a given impact parameter, the estimated for the USMgii absorbers tend to be higher compared to the Mg ii absorbers with low . These correlations imply that (i) at a given impact parameter, the absorbers with large tend to have large and (ii) for a given galaxies at higher impact parameters tend to have higher values of . Therefore, if GOTOQs follow this trend, we expect the for the host galaxies to be typically less than what has been found for USMgii host galaxies at high impact parameters. Indeed, Guha & Srianand (2023) have measured over the redshift range . Here, we check whether such measurements for individual cases in the case of GOTOQs follow the above-mentioned trend.
Accurate photometric measurements are required in several bands to measure the stellar masses () of GOTOQs using SED fitting techniques. This is possible only for 12 of the 39 objects with clear photometric extensions in our sample. For these twelve systems, the angular separations between the quasars and the foregrounds are large, so reliable fluxes in different photometric bands can be measured by subtracting the quasar PSF. We computed the foreground galaxy fluxes in the grz bands of the DECaLS photometry. Then, we fitted the spectral energy distributions (SED) using the publicly available code BAGPIPES (Carnall et al., 2018). During the fit, we assumed a delayed star formation history (Carnall et al., 2019) and kept the redshift constant as obtained from the [O ii] emission line. Obtained stellar masses of the GOTOQ with photometric counterparts are given in column 3 of Table 4. In Figure 9, we show plot vs. D for these objects and also show the best fit obtained for the host galaxies of USMgii and that obtained for the MAGIICAT galaxies. However, it is important to note that the correlation observed for the GOTOQs may be influenced by the mass-size relation of galaxies (Ward et al., 2024). Since more massive galaxies tend to be larger, massive GOTOQs are also larger in size and, consequently, may have larger impact parameters. It will be good to measure the host galaxy properties for more GOTOQs using HST photometry or AO-assisted ground-based imaging. This will provide more insights into how stellar mass controls the equivalent width distributions at low impact parameters.